### v2

@@ 10 Conclusions
  • Battery technology is evolving rapidly, with advancements in energy density, safety, and cost-effectiveness expected in the near future.
  • Higher voltage systems (700–800V) are gaining adoption, improving efficiency, reducing losses, and enabling faster charging, though they require updates in infrastructure and power electronics.
  • Sodium-based batteries are emerging as an alternative to lithium-ion batteries, offering advantages such as abundant raw materials, lower costs, and reduced environmental impact, though they remain in early development.
@@
  • TNO battery technology is gaining attention, offering high charging and discharging power capabilities, which could significantly reduce EV charging times and improve overall vehicle performance.
  • Battery swapping systems, such as those introduced by NIO, provide a promising alternative to fast charging, enabling quick battery replacements and addressing infrastructure limitations.
@@
  • The role of BMS is expanding, incorporating AI, advanced diagnostics, and cybersecurity measures to enhance performance and longevity.
  • Safety remains a critical priority, with stricter regulatory standards governing battery pack design, testing, and operational safety, particularly for thermal runaway prevention and high-voltage handling.
  • Thermal management is essential for high-performance battery systems, with ongoing advancements in cooling solutions for both fast charging and high-power applications.
  • Charging infrastructure must evolve alongside battery technology, as many existing chargers cannot accommodate 800V systems, requiring innovative solutions for compatibility with lower-voltage charging networks.
  • Second-life applications for EV batteries are expanding, including energy storage for the grid, supporting a more sustainable lifecycle and reducing environmental impact.
  • Battery electrification is extending beyond passenger cars, with significant research efforts in heavy-duty transport, buses, trucks, and aviation, each requiring unique design considerations.
  • Aviation electrification presents new challenges, requiring alternative battery architectures, extreme environmental durability, and cost-effective solutions distinct from ground vehicles.
  • The push for sustainable, cost-effective, and high-performance battery solutions will continue driving advancements in materials, design, and infrastructure, shaping the future of electric mobility.
@@ 11 List of abbreviations
    BTMS  Battery Thermal Management System.
@@
    ECE     Economic Commission for Europe.
@@
    EV        Electric Vehicle.
@@
    FEMA   Failure Mode and Effect Analysis.
@@
    HARA   Hazard Analysis and Risk Assessment.
@@
    HVIL   High-Voltage Interlock Loop.
@@
    ISO     International Organization for Standardization.
@@
    LCO    Lithium Cobalt Oxide.
@@
    MSD    Manual Service Disconnect.
@@
    NMC    Lithium Nickel Manganese Cobalt Oxide.

### v3

@@ 10 Conclusion
  • Battery technology is developing and advancements in energy density, safety, and cost-effectiveness are expected shortly.
  • High voltage systems (700–800 V) are developing due to their high efficiency, reduced power losses, and faster charging. Despite this, they require upgrades in infrastructure and power electronics.
  • Sodium-based batteries are developing as a great alternative to Li-ion batteries, offering advantages such as abundant raw materials, lower costs, and reduced environmental impact, while they remain in early development.
@@
  • TNO battery technology offers high charging and discharging power capabilities, which could significantly reduce EV charging times and improve overall the EV performance.
  • Battery swapping systems, such as those introduced by NIO, provide a great alternative to fast charging, enabling quick battery swaps and solving infrastructure limitations.
@@
  • The role of BMS is expanding lifetime, advanced diagnostics, and cybersecurity measures to enhance performance and longevity of the battery pack.
  • Safety is a critical priority, with stricter regulatory standards governing battery pack design, testing, and operational safety, particularly for thermal runaway prevention and high-voltage handling.
  • Thermal management is essential for high-performance BMS, with ongoing progress in cooling solutions for both fast charging and high-power applications.
  • Charging infrastructure must evolve as well as battery technology, because many existing chargers cannot accommodate 800 V systems, requiring new solutions for compatibility with lower-voltage charging networks.
  • Second-life applications for EV batteries are expanding, which includes energy storage for the grid and supporting a more sustainable lifecycle, and reducing environmental impact.
  • Battery electrification is extending beyond passenger cars, with significant research efforts in heavy-duty transport, such as buses, trucks, and aviation, each of which needs unique design considerations.
  • Aviation electrification have new challenges, requiring alternative battery architectures, enhanced environmental durability, and cost-effective solutions that differ from those used in ground vehicles (buses, cars, etc).
  • The sustainable development, cost-effective, and high-performance battery solutions will continue to propel advancements in materials, design, and infrastructure, defining the future of electric mobility.
@@ 11 List of abbreviations
    ECE     Economic Commission for Europe.
@@
    ECU     Electronic Control Unit.
@@
    FEMA   Failure Mode and Effect Analysis.
@@
    HARA   Hazard Analysis and Risk Assessment.
@@
    HVIL   High-Voltage Interlock Loop.
@@
    ISO     International Organization for Standardization.
@@
    LCO    Lithium Cobalt Oxide.
@@
    LFP    Lithium Iron Phosphate.
@@
    NMC    Lithium Nickel Manganese Cobalt Oxide.
@@
    SG       Safety Goal.